OpenStudy (anonymous):

What is the resulting ordered pair if the value of the independent variable is –3? f(x)=-2x^3-4x^2-1 A. (–17, –3) B. (17, –3) C. (–3, –91) D. (–3, 17)

OpenStudy (jdoe0001):

\(\bf f(x)=-2x^3-4x^2-1\qquad \textit{independent variable}=x=-3\\ \quad \\ f({\color{red}{ -3}})=-2({\color{red}{ -3}})^3-4({\color{red}{ -3}})^2-1\)

OpenStudy (anonymous):

I did that but i might ave gotten my math wrong i got 17

OpenStudy (anonymous):

Wait so its D?

OpenStudy (jdoe0001):

hmmm ok ... \(\bf f({\color{red}{ -3}})=-2({\color{red}{ -3}})^3-4({\color{red}{ -3}})^2-1\\ \quad \\ (-3)^3\implies -3\cdot -3\cdot -3\implies -27\\ (-3)^2\implies -3\cdot -3\implies +9\\ \quad \\ f({\color{red}{ -3}})=-2({\color{red}{ -27}})^3-4({\color{red}{ 9}})^2-1\)

OpenStudy (jdoe0001):

hmmm ahemm kind forgot to remove the exponents... one sec \(\bf f({\color{red}{ -3}})=-2({\color{red}{ -3}})^3-4({\color{red}{ -3}})^2-1\\ \quad \\ (-3)^3\implies -3\cdot -3\cdot -3\implies -27\\ (-3)^2\implies -3\cdot -3\implies +9\\ \quad \\ f({\color{red}{ -3}})=-2({\color{red}{ -27}})-4({\color{red}{ 9}})-1\)

OpenStudy (anonymous):

-91?

OpenStudy (jdoe0001):

actuially, I kinda used +2 my bad.. yes your'e correct is 17, so it's D

OpenStudy (anonymous):

Okay lol thanks